# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ 5 Experiments using simulated data
In the following, we conduct several experiments using simulated data. We compare the different preconditioners and methods for calculating predictive variances. Furthermore, we analyze our proposed methods concerning the accuracy and runtime for parameter estimation and prediction. Krylov subspace methods are compared to calculations based on the Cholesky decomposition using the GPBoost library. Note, however, that the Cholesky decomposition is also not exact since round-off errors can accumulate when using finite precision arithmetic. Additionally, we do a comparison with the lme4 and glmmTMB software libraries, which both rely on the Cholesky decomposition. Code to reproduce the simulated and the real-world data experiments of this article is available at https://github.com/pkuendig/KrylovGMMs.
@@ 5.1 Experimental setting
We simulate data from a model with two randomly crossed random effects. Unless stated otherwise, we use a balanced random effects design, in which both random effects have the same dimension $m_{1}=m_{2}$, and the random effect variances are $\sigma_{1}^{2}=\sigma_{2}^{2}=0.25$. But we also conduct experiments for an unbalanced design and other choices of random effects variances. We include a linear regression fixed effects term with five covariates plus an intercept for simulation, estimation, and prediction. The covariates are sampled from a normal distribution with mean zero and variance chosen such that the signal-to-noise ratio between the fixed and random effects is one, and the true regression coefficients are all one except for the intercept, which is zero. The response variable $y$ follows either a Gaussian likelihood with variance $\sigma^{2}=0.25$ or a Bernoulli likelihood with a logit link function. More details, e.g., on the sample size $n$ and the dimension $m$ of the random effects, are provided in the subsections below.
@@
For analyzing prediction accuracy, the data is randomly split into training and test data sets. We measure both the accuracy of point and probabilistic predictions for the test data random effects $\tilde{b}_{p}=Z_{po}b+Z_{pp}b_{p}$; see (8). Point predictions in the form of predictive means are evaluated using the root mean squared error (RMSE), and for probabilistic predictions, we use the log score (LS) $-\frac{1}{n_{p}}\sum_{i=1}^{n_{p}}\log(\mathcal{N}((\tilde{b}_{p})_{i};\omega_{p,i},(\Omega_{p})_{ii}))$, where $\omega_{p,i}$ and $(\Omega_{p})_{ii}$ are the predictive latent means and variances for the random effect $(\tilde{b}_{p})_{i}$ of test observation $i$, and $\mathcal{N}((\tilde{b}_{p})_{i};\omega_{p,i},(\Omega_{p})_{ii})$ denotes a Gaussian density evaluated at $(\tilde{b}_{p})_{i}$. We also measure the runtime in seconds.
@@
All calculations are done on a laptop with an Intel i7-12800H processor and 32 GB of random-access memory. We use the GPBoost library version 1.5.8 for Krylov subspace and Cholesky decomposition-based methods denoted as ‘Krylov (GPBoost)’ and ‘Cholesky (GPBoost)’, respectively. Cholesky-based calculations are additionally done with the R packages lme4 version 1.1-35.5 and glmmTMB version 1.1.10. For Gaussian likelihoods, parameter estimation is done by minimizing the negative log-marginal likelihood. For non-Gaussian likelihoods, all implementations use the Laplace approximation. For parameter estimation, we set the initial values for all models to the default values used by GPBoost. Besides the above-mentioned points, the default options of the packages are used, including package-specific internal optimizers.
@@
For the Krylov subspace methods, if not stated otherwise, we use the SSOR preconditioner, set the number of random vectors for STE and SLQ to $t=50$, and use a rank of $k=50$ for the two low-rank preconditioners. For the CG algorithm, we use a convergence tolerance of $10^{-2}$ for calculating marginal likelihoods and gradients, and a tolerance of $10^{-3}$ for predictive variances. Unless stated otherwise, we use the Algorithm 1 and $s=1000$ samples for calculating predictive variances. Further, we adopt a sample average approximation approach (Kim et al., 2015) when maximizing the marginal likelihood.
@@ 5.2 Preconditioner comparison
First, we compare the proposed preconditioners with regard to the accuracy and runtime for approximating log-marginal likelihoods. We set the total number of group levels to $m=100^{\prime}000$, use a sample size of $n=1^{\prime}000^{\prime}000$, and consider varying numbers of random vectors $t$ for the SLQ method. Note that these dimensions are approximately the largest ones such that Cholesky-based computations can be run in a “reasonable” amount of time which is roughly four hours for one likelihood evaluation. For each preconditioner and number of random vectors, the calculation of the likelihood is done at the true covariance parameters and repeated $100$ times with different random vectors. Figure 2 reports the results when the response variable follows a Gaussian likelihood. The runtime reported is the average wall-clock time. We first observe that the SSOR and ZIC preconditioners yield very accurate log-likelihood approximations with much smaller variances compared to the diagonal, Lanczos, and pivoted Cholesky preconditioners. This observation is consistent with the theoretical results presented in Section 4. Furthermore, we find that computations based on Krylov subspace methods are much faster than Cholesky-based computations. For instance, when using $t=50$ random vectors, we obtain a speed-up of approximately four orders of magnitude. Moreover, the runtimes of the CG method are similar for all preconditioners and even without preconditioning. In line with our convergence theory, the diagonal preconditioner has slightly longer runtimes compared to the SSOR preconditioner, but the differences are small. We conjecture that the finding that the preconditioned and non-preconditioned CG methods have similar runtimes can, partially, be explained by computational overheads of the preconditioners and other computations. The longest runtimes are observed for the Lanczos preconditioner. In summary, while preconditioning does not reduce the runtime, it can substantially reduce the variance of log-marginal likelihoods. In particular, variance reduction is very effective with the SSOR and ZIC preconditioner but has no clear effect with the diagonal, Lanczos, and pivoted Cholesky preconditioners.
Figure 2: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when the response variable follows a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
In Figure 9 in Appendix A.8, we report analogous results when the response variable follows a Bernoulli likelihood with a logit link function and all other settings are identical for the SSOR, ZIC, and diagonal preconditioners. Similar observations can be made in this non-Gaussian setting, with the SSOR and ZIC preconditioners having substantially smaller variances than the diagonal preconditioner. Furthermore, in Figure 10 in Appendix A.8, we report the results when using an unbalanced random-effects design with $m_{2}=\frac{m_{1}}{2}$, and in Figure 11 in Appendix A.8 we present the results when having a different signal-to-noise ratio with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$ for a Gaussian likelihood. The results for these two settings are again very similar to the previous ones. Cholesky-based computations are faster for the unbalanced design, since $Z^{T}WZ$ is more sparse, but still three orders of magnitude slower than Krylov subspace methods. Note that the experiments in this section and the following Section 5.3 are conducted on only one simulated data set since we do not want to mix sampling variability and variability of the simulation-based approximations. However, the results do not change when using other samples (results not tabulated).
@@ 5.3 Comparing methods for predictive variances
Next, we compare the accuracy of the different approximations for predictive variances introduced in Section 3.2 using simulated data with $n=n_{p}=100^{\prime}000$ training and test points and a total of $m+m_{p}=10^{\prime}000$ group levels. The response variable follows a Gaussian likelihood, and we calculate the RMSE of the predictive variances compared to the “exact” Cholesky-based results. We also measure the runtime for prediction, which includes the calculation of the latent predictive means and variances. Predictive distributions are calculated at the true covariance parameters.
Figure 3 shows the RMSE versus the wall-clock time for different numbers of random vectors $s$ for the stochastic methods and different ranks $k$ for the Lanczos-based approximation. For the former, we average the results over $100$ independent repetitions for every $s$ and add whiskers to the plot representing confidence intervals for the RMSE obtained as ±2$\pm$ 2$\times$ 2 × standard errors, but the corresponding whiskers are not visible since they are very small. By far the most accurate predictive variances for a given runtime are obtained with the simulation-based Algorithm 1. For instance, for $s=200$, predictive variances are very accurate with an RMSE of approximately $2.6\times 10^{-4}$, and the runtime is more than $100$ times faster compared to Cholesky-based calculations. Algorithm 2 achieves the second highest prediction accuracy for a given runtime, and Algorithm 3 is considerably slower and less accurate. Finally, Lanczos-based approximations are very inaccurate and the RMSE decreases only very slowly with increasing rank $k$.
Figure 3: Comparison of different methods for predictive variances. The number of random vectors $s$ and the Lanczos rank $k$ are annotated in the plot.
@@ 5.4 Accuracy of parameter estimation and predictive distributions
In the following, we analyze the properties of the variance parameter and linear regression coefficient estimators and the accuracy of predictive distributions for Gaussian and Bernoulli likelihoods. We set the total number of group levels to $m+m_{p}=4^{\prime}000$ and simulate $n=n_{p}=40^{\prime}000$ training and test points. This is roughly the largest number of group levels such that we can run calculations with the glmmTMB and lme4 packages in a reasonable amount of time. We perform $100$ simulation repetitions, and prediction is done using estimated parameters. For Bernoulli likelihoods, repeating the estimation $100$ times with lme4 leads to very long runtimes; see also Section 5.5. Therefore, we do not report estimates and prediction accuracy for this case.
Figure 4 visualized the estimates for the variance parameters. The RMSE and bias of the variance and coefficient estimators are additionally reported in Appendix A.9 in Tables 2 and 3 for Gaussian and Bernoulli likelihoods, respectively. We observe virtually no differences among the estimates obtained using Krylov subspace methods and Cholesky-based computations with the GPBoost, lme4, and glmmTMB packages. Note that the variance parameter estimators for Bernoulli likelihoods are slightly downward biased due to the Laplace approximation and the moderate sample size.
Figure 4: Estimated variance parameters. The red rhombi represent means. The dashed lines indicate the true values. For the Bernoulli likelihood the estimates for lme4 are not computed due to excessively long runtimes.
@@
5.5 Runtime comparison
In this subsection, we compare the runtime for parameter estimation when using Krylov subspace methods and Cholesky-based calculations with the GPBoost, lme4, and glmmTMB packages. For this, we consider different dimensions of the random effects ranging from $m=1^{\prime}000$ up to $m=1^{\prime}000^{\prime}000$ and $n=10m$ using the same setting as described in Section 5.1. Figure 6 shows the wall-clock time in seconds versus the total number of random effects $m$. Krylov subspace methods achieve the fastest runtime for all $m$. For instance, for $m=20^{\prime}000$ and a Gaussian likelihood, Krylov subspace methods are approximately three orders of magnitudes faster than Cholesky-based computations when using the GPBoost package. For a Gaussian likelihood, glmmTMB is approximately four orders of magnitudes slower than the Krylov subspace methods implemented in GPBoost, and lme4 is approximately four orders of magnitudes slower than the Krylov subspace methods for a Bernoulli likelihood. In general, estimation takes the longest with glmmTMB for Gaussian likelihoods, and for Bernoulli likelihoods, estimation takes the longest with lme4.
In Figure 12 in Appendix A.9, we additionally compare the runtime for parameter estimation for an unbalanced random effects design with $m_{2}=\frac{m_{1}}{2}$. Cholesky-based computations are faster compared to the balanced design due to the increased sparsity in $Z^{T}WZ$, but Krylov subspace methods again clearly achieve the fastest runtime for all $m$. For instance, they are three orders of magnitude faster than Cholesky-based computations using GPBoost for $m=50^{\prime}000$ and a Gaussian likelihood. Note that it would also be interesting to perform a runtime comparison for calculating log-marginal likelihoods as this would eliminate the impact of using different optimizers across different libraries, but the lme4 and glmmTMB packages do not support the explicit calculation of marginal likelihoods. However, ‘Krylov (GPBoost)’ and ‘Cholesky (GPBoost)’ differ only in the use of Krylov subspace methods vs. Cholesky-based computations, and this comparison therefore clearly demonstrates the speed-up with Krylov subspace methods.

### v2

@@ 5 Experiments using simulated data
In the following, we conduct experiments using simulated data. We compare the different preconditioners and methods for calculating predictive variances. Furthermore, we analyze our proposed methods concerning the accuracy and runtime for parameter estimation and prediction. Krylov subspace methods are compared to computations relying on sparse Cholesky decompositions. Note, however, that Cholesky decompositions also do not necessarily yield exact results since round-off errors can accumulate when using finite precision arithmetic. Code to reproduce the simulated and the real-world data experiments of this article is available at https://github.com/pkuendig/KrylovGMMs.
@@ 5.1 Experimental setting
We simulate data from a model with two crossed random effect components. Unless stated otherwise, we use a balanced random effects design, i.e., each random effect $b_{k,j}$, $j=1,\dots,m_{k}$, $k=1,2$, occurs the same number of times, $\sum_{i=1}^{n}(Z_{1})_{ij}=\sum_{i=1}^{n}(Z_{2})_{ij}=n/m_{1}=n/m_{2}$, both random effect components $b_{1}$ and $b_{2}$ have the same dimension $m_{1}=m_{2}$, the crossing structure between the two random effect components is random, and the random effects variances are $\sigma_{1}^{2}=\sigma_{2}^{2}=0.25$. A random crossing structure is generated by generating incidence matrices $Z_{k}\in\{0,1\}^{n\times m_{k}}$ as follows. We first assign exactly $n/m_{k}$ ones to column $j$ in consecutive rows, $(Z_{k})_{ij}=1\quad\text{for }i=(j-1)n/m_{k}+1,\dots,jn/m_{k}$, $j=1,\dots,m_{k}$, and zero otherwise, and we randomly permute the rows of $Z_{2}$. We consider various sample sizes $n$ and dimensions $m$ of the random effects as described in the subsections below. The response variable $y$ follows either a Gaussian likelihood with error variance $\sigma^{2}=0.25$ or a Bernoulli likelihood with a logit link function. We additionally conduct experiments for less regular unbalanced designs, unequal dimensions of the random effect components ($m_{1}\neq m_{2}$), and other signal-to-noise ratios. See the subsections below for more information. For the experiments regarding the accuracy and runtimes of parameter estimation and prediction in Sections 5.5 and 5.4, we include a fixed effects linear predictor term with five covariates and an intercept. The regression coefficients are all set to one except for the intercept, which is set to zero, and the covariates are sampled from a normal distribution with mean zero and variance chosen such that the fixed and random effects have equal variance.
@@
For analyzing prediction accuracy, the data is randomly split into training and test data sets. We measure both the accuracy of point and probabilistic predictions for the test random effects $\tilde{b}_{p}=Z_{po}b+Z_{pp}b_{p}$; see (8). Point predictions in the form of predictive means are evaluated using the root mean squared error (RMSE), and for probabilistic predictions, we use the log score (LS) $\frac{1}{n_{p}}\sum_{i=1}^{n_{p}}\log(\mathcal{N}((\tilde{b}_{p})_{i};\omega_{p,i},(\Omega_{p})_{ii}))$, where $\omega_{p,i}$ and $(\Omega_{p})_{ii}$ are the predictive latent means and variances for the random effect $(\tilde{b}_{p})_{i}$ of test observation $i$, and $\mathcal{N}((\tilde{b}_{p})_{i};\omega_{p,i},(\Omega_{p})_{ii})$ denotes a Gaussian density evaluated at $(\tilde{b}_{p})_{i}$. We also measure the runtime in seconds.
@@
All calculations are done on a laptop with an Intel i7-12800H processor and 32 GB of random-access memory. We use the GPBoost library version 1.6.4 for Krylov subspace and Cholesky decomposition-based methods denoted as ‘Krylov (GPBoost)’ and ‘Cholesky (GPBoost)’, respectively. Cholesky-based calculations are additionally done with the R packages lme4 version 1.1-35.5 and glmmTMB version 1.1.10. For Gaussian likelihoods, parameter estimation is done by minimizing the negative log-marginal likelihood. For non-Gaussian likelihoods, all implementations use the Laplace approximation. For parameter estimation, we set the initial values for all models to the default values used by GPBoost. Besides the above-mentioned points, the default options of the packages are used, including package-specific internal optimizers.
@@
For the Krylov subspace methods, if not stated otherwise, we use the SSOR preconditioner and $t=50$ random vectors for STE and SLQ. For the CG algorithm, we use a convergence tolerance of $10^{-2}$ for calculating marginal likelihoods and gradients and a tolerance of $10^{-3}$ for predictive variances. Unless stated otherwise, we use Algorithm 1 and $s=1000$ samples for calculating predictive variances. Further, we adopt a sample average approximation approach (kim2015guide) when maximizing the marginal likelihood.
@@ 5.2 Preconditioner comparison
We compare the preconditioners introduced in Section 3.1 with regard to the accuracy and runtimes for approximating log-marginal likelihoods. We use a total number of $m=100,000$ random effects and a sample size of $n=1,000,000$. These dimensions are approximately the largest such that Cholesky-based computations can be run in a reasonable amount of time (approx. four hours for one likelihood evaluation). Furthermore, we use both a balanced random effects design as described in Section 5.1 and a less regular unbalanced design. The latter is generated as described in Appendix A.7. The calculation of the marginal likelihoods is done at the data-generating parameters and repeated $100$ times with different random vectors for the SLQ method.
First, we analyze how the accuracy and runtimes depend on the number of random vectors $t$ in the SLQ method for a Gaussian likelihood. Figure 1 reports standard deviations of the stochastic estimates and average wall-clock times for the different preconditioners for both the balanced and unbalanced random effects designs. The figure also reports the number of iterations (bottom row, right axis) of the CG method used for calculating the quadratic from in the log-marginal likelihood. The wall-clock time and negative log-marginal likelihood when using a Cholesky decomposition are annotated in the plot. We first observe that the SSOR and ZIC preconditioners yield the most accurate log-marginal likelihood approximations. Compared to the diagonal preconditioner and the unpreconditioned SLQ method, the SSOR and ZIC preconditioners yield standard deviations that are approximately one order of magnitude lower for both random effects designs. This observation is consistent with the theoretical results presented in Section 4. Furthermore, as expected from theory, the diagonal preconditioner does not give any variance reduction for the balanced design, but for the unbalanced design, we observe smaller standard deviations compared to the unpreconditioned CG method. The runtimes and the number of CG iterations are similar for all preconditioners and also without preconditioning for the balanced design. In contrast, for the unbalanced design, all preconditioners substantially reduce the runtimes and numbers of CG iterations. In line with the convergence theory in Section 4, the diagonal preconditioner has a slightly higher number of CG iterations and runtimes compared to the SSOR preconditioner, but the differences are small. Note that the SSOR preconditioner has a higher “overhead” per CG iteration than the diagonal preconditioner since additional sparse triangular solves are done for the SSOR preconditioner. Finally, we observe that the Krylov subspace-based methods are much faster than Cholesky-based calculations. For instance, when using $t=50$ random vectors, they result in a speedup of approximately four orders of magnitude. Note that all preconditioners yield unbiased SLQ-approximated log-marginal likelihoods. To empirically verify this, we additionally report in Figure 8 in Appendix A.8 histograms of log-marginal likelihoods.
Figure 1: Accuracy (in standard deviations; top row), wall-clock time (in seconds; bottom row, left axis), and number of CG iterations (bottom row, right axis) for different preconditioners and varying numbers of random vectors $t$ in the SLQ method for calculating log-marginal likelihoods for a Gaussian likelihood and balanced and unbalanced random effects designs.
In Figures 9 and 10 in Appendix A.8, we report analogous results when the likelihood is Bernoulli with a logit link function and all other settings are identical. We find qualitatively very similar results as in the Gaussian setting. In particular, Krylov subspace-based methods are approximately $50,000$ times faster than Cholesky-based computations. The SSOR and ZIC preconditioners yield substantially more accurate log-marginal likelihood approximations than the diagonal preconditioner and the unpreconditioned case. Next, in Figure 11 in Appendix A.8, we present additional results when having a higher signal-to-noise ratio obtained by using a lower error variance $\sigma^{2}=0.05$ for a Gaussian likelihood. The results are again qualitatively similar to the ones of the lower signal-to-noise ratio. We further analyze the impact of the irregularity of the random effects design. For this, we vary the size parameter $r$ in the unbalanced design generation approach explained in Appendix A.7. The results are reported in the left column of Figure 12 in Appendix A.8. We find that the runtime of the unpreconditioned CG method increases with increasing irregularity, whereas the runtimes remain constant for the preconditioned CG method for all preconditioners. The accuracy of the SLQ method is only marginally affected by the irregularity of the design with and without preconditioning. In the right column of Figure 12 in Appendix A.8, we also report results when varying the random effects dimension $m$. An unbalanced design and a constant ratio $n/m=10$ (as above) is used for this experiment. We find that there is a small increase in the variance of the SLQ method, the numbers of CG iterations are essentially constant, but the runtimes of the CG method increases due to a higher computational complexity per CG iteration when $m$ increases.
@@ 5.3 Comparing methods for predictive variances
Next, we compare the accuracy of the different methods for calculating predictive variances, as introduced in Section 3.2, using simulated data with $n=n_{p}=100,000$ training and test points and a total of $m+m_{p}=10,000$ group levels. The response variable follows a Gaussian likelihood, and we calculate the RMSE of the approximate predictive variances to the “exact” Cholesky-based predictive variances. We also measure the runtime for prediction, which includes the calculation of the latent predictive means and variances. Predictive distributions are calculated at the data-generating variance parameters.
Figure 2 visualizes the accuracy (RMSE) versus the wall-clock time for different numbers of random vectors $s$ for the stochastic methods and different ranks $k$ for the Lanczos-based approximation. For the former, we average the results over $100$ independent repetitions for every $s$ and add confidence intervals for the RMSE obtained as $\pm 2\times$ standard errors (not visible since they are very small). The runtime of Cholesky-based calculations is annotated in the plot. By far the most accurate predictive variances for a given runtime are obtained with the simulation-based Algorithm 1. For instance, for $s=200$, predictive variances are very accurate with an RMSE of approximately $2.6\times 10^{-4}$, and the runtime is more than $100$ times faster compared to Cholesky-based calculations. Algorithm 2 achieves the second highest prediction accuracy for a given runtime, and Algorithm 3 is considerably slower and less accurate. Finally, Lanczos-based approximations are very inaccurate and the RMSE decreases only very slowly with increasing rank $k$.
Figure 2: Accuracy-runtime comparison of different methods for predictive variances. The number of random vectors $s$ and the Lanczos rank $k$ are annotated in the plot.
5.4 Runtime comparison for parameter estimation
In this subsection, we compare the runtime for parameter estimation when using Krylov subspace methods and Cholesky-based calculations with the GPBoost, lme4, and glmmTMB packages. We consider different dimensions of the random effects ranging from $m=1,000$ up to $m=1,000,000$ and $n=10m$ using the setting described in Section 5.1. In particular, we use two crossed random effect components and include five covariates plus an intercept in a linear predictor. Figure 3 shows the wall-clock time in seconds versus the number of random effects $m$. Krylov subspace methods achieve the fastest runtime for all $m$. For instance, for $m=20,000$ and a Gaussian likelihood, Krylov subspace methods are approximately three orders of magnitudes faster than Cholesky-based computations when using the GPBoost package. For a Gaussian likelihood, glmmTMB is approximately four orders of magnitudes slower than the Krylov subspace methods implemented in GPBoost, and lme4 is approximately four orders of magnitudes slower than the Krylov subspace methods for a Bernoulli likelihood. In general, estimation takes the longest with glmmTMB for Gaussian likelihoods, and for Bernoulli likelihoods, estimation takes the longest with lme4.
In Figure 13 in Appendix A.9, we additionally compare the runtime for parameter estimation for a random effects design with $m_{2}=2$. Cholesky-based computations are faster compared to the case $m_{1}=m_{2}$ due to the increased sparsity in $Z^{T}WZ$, but Krylov subspace methods again clearly achieve the fastest runtime for all $m$. For instance, they are three orders of magnitude faster than Cholesky-based computations when using GPBoost for $m=50,000$ and a Gaussian likelihood.
@@ 5.5 Accuracy of parameter estimation and predictive distributions
In the following, we analyze the properties of the variance parameter and linear regression coefficient estimates and the accuracy of predictive distributions for Gaussian and Bernoulli likelihoods when using Krylov-subspace methods. As mentioned in Section 5.1, we include five covariates plus an intercept in a linear predictor. We set the total number of training and prediction random effects to $m+m_{p}=4,000$ and simulate $n=n_{p}=40,000$ training and test points. This is roughly the largest number of group levels such that we can run calculations with the glmmTMB and lme4 packages in a reasonable amount of time. We perform $100$ simulation repetitions, and prediction is done using estimated parameters. For Bernoulli likelihoods, repeating the estimation $100$ times with lme4 leads to very long runtimes; see also Section 5.4. Therefore, we do not report estimates and prediction accuracy for this case.
Figure 4 visualizes the estimates for the variance parameters. The RMSE and bias of the variance and regression coefficient estimates are additionally reported in Appendix A.9 in Tables 2 and 3 for Gaussian and Bernoulli likelihoods, respectively. We observe virtually no differences among the estimates obtained using Krylov subspace methods and Cholesky-based computations with the GPBoost, lme4, and glmmTMB packages. Note that the variance parameter estimates for Bernoulli likelihoods are slightly downward biased due to the Laplace approximation and the moderate sample size.
Figure 4: Estimated variance parameters. The red rhombi represent means. The dashed lines indicate the true values. For the Bernoulli likelihood, the estimates for lme4 are not computed due to excessively long runtimes.